# Prosoluble subgroups of the profinite completion of the fundamental group of compact 33 33-manifolds

### v2

@@ 5 The finitely generated case
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group. If $H$ is a finitely generated prosoluble subgroup of $G$, then $H$ embeds into a free profinite product $\coprod_{v\in\Gamma}H_{v}$ of finitely many finitely generated profinite groups $H_{v}=H\cap\mathcal{G}(v)^{g}$ and $g\in G$. Moreover, if $H\neq H_{v}$ for some $v$, then either $H_{v}$ are finitely generated pro-$p$ or $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group (i.e., $C$ is cyclic, $|C|$ divides $p-1$ for any $p\in\pi$ and $[c,z]\neq 1$ for all $1\neq c\in C,0\neq z\in\widehat{\mathbb{Z}}_{\pi}$).
@@
Note that the abstract subgraph $D=\{m\mid H_{m}\neq 1\}$ has finite diameter by Theorem 3.5, and by Proposition 4.10 $D$ has finitely many connected components up to translation. Therefore $H\backslash D$ has finitely many connected components and so by Corollary 3.7 there exists a simply connected $H$-quotient graph $\overline{S}$ on which $H$ acts with trivial edge stabilizers.
By item (i) of Theorem 2.23, $(H,V(\overline{S}))$ is a $\mathcal{S}$-projective pile where $\mathcal{S}$ is the class of all finite soluble groups. By [Zal24, Remark 4.9] there exists a second countable space $T$ with the same non-trivial point stabilizers such that $(H,T)$ is a projective pile. Since $T$ is second countable, the quotient map $T\to H\backslash T$ has a continuous section (see [RZ10, Lemma 5.6.7] or [Rib17, 1.3.10]). By Theorem 2.23 $(H,T)$ is a $\mathcal{S}$-projective pile such that the quotient map $T\to H\backslash T$ has a continuous section. Therefore by Proposition 2.24 $H$ embeds into a free prosoluble product
@@
where $F$ is a free prosoluble group. Moreover, every $H_{t}$ is of the form $H\cap G(v)^{g}$. By Corollary 4.9 $H_{t}$ are either all pro-$p$ or $H$ is Frobenius and all $H_{t}$ are conjugate as they are Hall subgroups of $H$. Then by [Zal24, Theorem 1.4] $H$ embeds into a desired free profinite product. ∎
@@ 6 Hyperbolic virtually special groups
A group is virtually special if there is a finite index subgroup isomorphic to the fundamental group of a compact special cube complex. A group is virtually compact special if there is a finite index subgroup isomorphic to the fundamental group of a compact special cube complex [HW08, Definition 3.2].
@@
A hierarchy of groups of length 00 0 is a single vertex labeled by a group. A hierarchy of groups of length $n$ is a graph of groups $(\mathcal{G}_{n},\Gamma_{n})$ together with hierarchies of length $n-1$ on each vertex of $\Gamma_{n}$ (cf. [WZ17, Definition 1.3]).
In [WZ17, Theorem 1.4] it is shown that a hyperbolic group $G$ is virtually compact special if, and only if, $G$ has a subgroup $\Gamma_{0}$ of finite index with a malnormal hierarchy. It is also shown that, in this case, $\Gamma_{0}$ has a 11 11-acylindrical action on a profinite tree (see [WZ17, Lemma 7.3]). In the relatively hyperbolic virtually compact special case, $\Gamma_{0}$, in addition, has a malnormal hierarchy terminating in parabolic subgroups (see [Zal23a, Theorem 3.2]).
@@
Let $G$ be a finitely generated hyperbolic virtually compact special group and $H$ a finitely generated prosoluble subgroup of $\widehat{G}$. Let $H_{0}=\widehat{\Gamma}_{0}\cap H$ where $\Gamma_{0}$ is the subgroup defined above. We use induction on the length of the malnormal hierarchy described above. Since $\widehat{\Gamma_{0}}$ is a fundamental group of a graph of profinite groups whose action on the standard graph is 11 11-acylindrical (see [WZ17, Lemma 7.3]), by Theorem F, we have an embedding
$H_{0}\hookrightarrow\coprod H_{0}\cap G(v)^{g}.$
Then by Corollary 4.3 either $H_{0}\leq H_{v}^{g}$ or all $H_{0}\cap G(v)^{g}$ are pro-$p$. In the first case the result follows immediately from the induction hypothesis. In the second case we observe first that, by Theorem F, $H_{0}\cap G(v)^{g}$ are finitely generated which allows to apply (cf. [WZ17, Theorem F]) to deduce that $H_{0}\cap G(v)^{g}$ are free pro-$p$. It follows that $H_{0}$ is projective and so $H$ is virtually projective. Since it is torsion-free, it is projective (see [Zal04, Proposition 2.5]). ∎
Theorem B yields a description of prosoluble subgroups of closed hyperbolic 33 33-manifolds.
@@
Let $\pi_{1}(M)$ be the fundamental group of a compact hyperbolic 33 33-manifold $M$ and $H$ a finitely generated prosoluble subgroup of the profinite completion $\pi_{1}(M)$. Then $H$ is projective.
@@
Let $G$ be a residually finite group hyperbolic relative to a malnormal family of parabolic subgroups $\mathcal{P}=\{P_{1},...,P_{n}\}$. A closed subgroup of the profinite completion of some $P_{i}$ will be called parabolic subgroup of $\widehat{G}$.
@@
    $H_{0}$ is a subgroup of a parabolic subgroup.
@@
Let $G$ be a finitely generated relatively hyperbolic virtually compact special group and $H$ a finitely generated prosoluble subgroup of $\widehat{G}$. In [Zal23a, Proof of Theorem 1.3] we see that there is a subgroup $\Gamma_{0}$ such that $\widehat{\Gamma_{0}}$ acts 11 11-acylindrically on a profinite tree. Let $H_{0}=\widehat{\Gamma}_{0}\cap H$. As in the proof of Theorem B, by Theorem F, we have an embedding
@@
In the second case we observe first that by Corollary 4.10 $H_{0}\cap G(v)^{g}$ are finitely generated which allows to apply (cf. [Zal23a, Theorem 1.4]) to deduce that $H_{0}\cap G(v)^{g}$ are free pro-$p$ products of parabolic pro-$p$ groups and a free pro-$p$ group. ∎
@@
Corollary 6.6.
Let $G$ be a torsion-free relatively hyperbolic virtually compact special group and $H$ a finitely generated prosoluble subgroup of the profinite completion $\widehat{G}$ of $G$. Then $H$ is virtually abelian or virtually a subgroup of a free prosoluble product of abelian pro-$p$ groups.
@@ 7 The fundamental group of standard compact arithmetic manifolds
It is shown in [BHW11] that standard uniform arithmetic subgroups of $\operatorname{SO}(n,1)$ are virtually compact special. Thus, Theorem B can be applied to standard arithmetic hyperbolic manifolds. Also, it is well-known that the fundamental group of these manifolds is hyperbolic relative to virtually abelian subgroups. Therefore, Proposition 6.4 yields
Theorem D.
Let $M$ be a standard uniform hyperbolic arithmetic manifold and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then $H$ is projective.
8 The fundamental group of compact 33 33-manifolds
@@
[WZ18, Theorem 1.3] A finitely generated pro-$p$ subgroup of the profinite completion of the fundamental group of a compact, orientable 33 33-manifold $M$ is a free pro-$p$ product of the pro-$p$ groups from the following list of isomorphism types:
@@
    For $p=2$: in addition to the list of (i) we have $C_{2^{m}}$; $D_{2^{k}}$; $Q_{2^{n}}$; $\mathbb{Z}_{2}\rtimes C_{2}$; the torsion-free extensions of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ by one of $C_{2}$, $C_{4}$, $C_{8}$, $D_{2}$, $D_{4}$, $D_{8}$, $Q_{16}$; the pro-22 22 extension of the Klein-bottle group $\mathbb{Z}\rtimes\mathbb{Z}$; the pro-22 22 completion of all torsion-free extensions of a soluble group $(\mathbb{Z}\rtimes\mathbb{Z})\rtimes\mathbb{Z}$ with a group of order at most 22 22.
@@
Let $M$ be a compact orientable 33 33-manifold. If $H$ is a finitely generated prosoluble subgroup of $\widehat{\pi_{1}(M)}$, then one of the following statements holds:
@@
    $H$ is isomorphic to a subgroup of a free prosoluble product of the pro-$p$ groups from the following list of isomorphism types:
@@
[…]
@@
    $H$ is a subgroup of a central extension of either $D_{2n}$ or tetrahedral group $T$ or octahedral group $O$ or a icosahedral group $I$ by a cyclic group of even order;
@@
    $H$ is a subgroup of the profinite completion of a 33 33-dimensional Bieberbach group $B$ (i.e, $B$ is torsion-free virtually $\mathbb{Z}^{3}$);
@@
    $H$ is a subgroup of the profinite completion of a group extension of $\mathbb{Z}^{2}\rtimes_{A}\mathbb{Z}$ by $C$ where $C$ is trivial or $C_{2}$ and $A$ is an Anosov matrix.
@@
    $H$ is an extension of a torsion-free procyclic group ^$\widehat{\mathbb{Z}}_{\sigma}$ by a subgroup $H_{0}$ of a finite free prosoluble product of finite cyclic $p$-groups with $H_{0}$ acting either trivially on ^$\widehat{\mathbb{Z}}_{\sigma}$ or by inversion.
@@ 8.1 Reduction to the irreducible case
Let $M$ be a compact, orientable, non-geometric 33 33-manifold. Since every compact manifold is a retract of a closed manifold, it follows easily that we may reduce to the case in which $M$ is closed.
@@
where each $M_{i}$ is prime 33 33-manifold. A prime 33 33-manifold is either irreducible or $S^{2}\times S^{1}$ or the non-orientable $S^{2}$ bundle over $S^{1}$. In the last two cases, the fundamental group is isomorphic to $\mathbb{Z}$ (see [Hem04] and [GSW21, Page 5]).
@@
Let $H$ be a finitely generated prosoluble subgroup of the profinite completion $G=\widehat{\pi_{1}(M)}$ of a reducible 33 33-manifold having non-trivial free product decomposition
@@
where each $M_{i}$ is irreducible 33 33-manifold and $F$ is a free group of finite rank. Then one of the following holds:
@@
    all intersection $H\cap\widehat{\pi_{1}(M_{i})}^{g}$, $g\in G$ are pro-$p$ for some fixed prime $p$ and $\sum_{v}d(H\cap\widehat{\pi_{1}(M_{i})}^{g})\leq d(H)$;
@@
    $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group (i.e $C$ is cyclic, $|C|$ divides $p-1$ for any $p\in\pi$ and $[c,z]\neq 1$ for all $1\neq c\in C,0\neq z\in\widehat{\mathbb{Z}}_{\pi}$);
@@ 8.2 Reduction to the geometric cases
Let $M$ be a closed irreducible 33 33-manifold. Then the action of $\widehat{\pi_{1}(M)}$ on the standard profinite tree $S$ associated to the JSJ-decomposition of $M$ is 44 44-acylindrical.
Then we can apply Theorem F to deduce
@@
Let $M$ be a closed 33 33-manifold. If $H$ is a finitely generated prosoluble subgroup of $\widehat{\pi_{1}(M)}$, then $H$ embeds into a free profinite product $\coprod_{v\in V}H_{v}$ of $H$-stabilizers of vertices in $S$ and one of the following statements holds:
@@
    $H$ is a profinite Frobenius group ^$\widehat{\mathbb{Z}}_{\pi}\rtimes C$;
@@
If $H$ satisfies $(i)$, then we can apply Theorem 8.1 again and deduce $(i)$ or $(ii)$ of Theorem E. The case $(ii)$ corresponds to Theorem E $(ii)$. Therefore we may assume that $H$ satisfies item $(iii)$ and since $H$ in this case is conjugate into a vertex of $\widehat{\pi_{1}M}$ we may assume that $H$ is contained in the profinite completion of a geometric manifold.

### v3

@@ 5 The finitely generated case
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite group and $G$ its fundamental group. If $H$ is a finitely generated prosoluble subgroup of $G$, then $H$ embeds into a free profinite product $\coprod_{v\in V(\Gamma)}H_{v}$ of finitely many finitely generated profinite groups $H_{v}=H\cap G(v)^{g}$ for some $g\in G$. Moreover, if $H\neq H_{v}$ for some $v$, then either $H_{v}$ is finitely generated pro-$p$ or $H\simeq\widehat{\mathbb{Z}}_{\pi}\rtimes C$ is a profinite Frobenius group.
@@
Note that the abstract subgraph $D=\{m\mid H_{m}\neq 1\}$ has finite diameter by Theorem 3.5, and by Proposition 4.10 $D$ has finitely many connected components up to translation. Therefore $\D\displaystyle H\backslash$ D has finitely many connected components and so by Corollary 3.7 there exists a simply connected $H$-quotient graph $\overline{S}$ on which $H$ acts with trivial edge stabilizers.
By item (i) of Theorem 2.24, $(H,V(\overline{S}))$ is an $\mathcal{S}$-projective group-pile where $\mathcal{S}$ is the class of all finite soluble groups. By [Zal24, Remark 4.9] there exists a second countable space $T$ with the same non-trivial point stabilizers such that $(H,T)$ is a projective group-pile. Since $T$ is second countable, the quotient map $\T\displaystyle T\to H\backslash$ T has a continuous section (see [RZ10, Lemma 5.6.7] or [Rib17, 1.3.10]). By Theorem 2.24 $(H,T)$ is an $\mathcal{S}$-projective group-pile such that the quotient map $\T\displaystyle T\to H\backslash$ T has a continuous section. Therefore by Proposition 2.25 $H$ embeds into a free prosoluble product
@@
where $F$ is a free prosoluble group. Moreover, every $H_{t}$ is of the form $H\cap G(v)^{g}$. By Corollary 4.9, the $H_{t}$ are either all pro-$p$, or $H$ is Frobenius and all $H_{t}$ are conjugate as they are Hall subgroups of $H$. Then by [Zal24, Theorem 1.4] $H$ embeds into a desired free profinite product. ∎
@@ 6 Hyperbolic virtually special groups
A group is virtually special if there is a finite index subgroup isomorphic to the fundamental group of a special cube complex. A group is virtually compact special if there is a finite index subgroup isomorphic to the fundamental group of a compact special cube complex whose hyperplanes satisfy some extra conditions [HW08, Definition 3.2].
@@
A hierarchy of groups of length $0$ is a single vertex labeled by a group. A hierarchy of groups of length $n$ is a graph of groups $(\mathcal{G}_{n},\Gamma_{n})$ together with hierarchies of length $n-1$ on each vertex of $\Gamma_{n}$ (cf. [WZ17, Definition 1.3]).
In [WZ17, Theorem 1.4] it is shown that a hyperbolic group $G$ is virtually compact special if, and only if, $G$ has a subgroup $\Gamma_{0}$ of finite index with a malnormal hierarchy. It is also shown that, in this case, $\Gamma_{0}$ has a $1$-acylindrical action on a profinite tree (see [WZ17, Lemma 7.3]). In the relatively hyperbolic virtually compact special case, $\Gamma_{0}$, in addition, has a malnormal hierarchy terminating in parabolic subgroups (see [Zal23a, Theorem 3.2]).
@@
Let $G$ be a finitely generated hyperbolic virtually compact special group and $H$ a finitely generated prosoluble subgroup of $\widehat{G}$. Let $H_{0}=\widehat{\Gamma}_{0}\cap H$ where $\Gamma_{0}$ is the subgroup defined above. We use induction on the length of the malnormal hierarchy described above. Since $\widehat{\Gamma_{0}}$ is a fundamental group of a graph of profinite groups $(\mathcal{G},\Gamma)$ whose action on the standard graph is $1$-acylindrical (see [WZ17, Lemma 7.3]), by Theorem F, we have an embedding
$H_{0}\hookrightarrow\coprod_{v\in V(\Gamma)}H_{0}\cap G(v)^{g_{i}},g_{i}\in\widehat{\Gamma}_{0}.$
Then by Corollary 4.3 either $H_{0}\leq H_{v}^{g}$ or all $H_{0}\cap G(v)^{g}$ are pro-$p$. In the first case the result follows immediately from the induction hypothesis. In the second case we observe first that, by Theorem F, $H_{0}\cap G(v)^{g}$ are finitely generated which allows us to deduce that $H_{0}\cap G(v)^{g}$ are free pro-$p$ (cf. [WZ17, Theorem F]). It follows that $H_{0}$ is projective and so $H$ is virtually projective. Since it is torsion-free, it is projective (see [Zal04, Proposition 2.5]). ∎
Theorem B yields a description of prosoluble subgroups of closed hyperbolic $3$-manifolds.
@@
Let $\pi_{1}(M)$ be the fundamental group of a closed hyperbolic $3$-manifold $M$ and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then $H$ is projective.
@@
Let $G$ be a residually finite group hyperbolic relative to a malnormal family of parabolic subgroups $\mathcal{P}=\{P_{1},...,P_{n}\}$. Any subgroup of a conjugate of $P_{i}$ will be called parabolic. A closed subgroup of the profinite completion of a maximal parabolic subgroup of $G$ be called parabolic subgroup of $\widehat{G}$.
@@
    $H_{0}$ is isomorphic to a subgroup of a parabolic subgroup.
@@
Let $G$ be a finitely generated relatively hyperbolic virtually compact special group and $H$ a finitely generated prosoluble subgroup of $\widehat{G}$. In [Zal23a, Proof of Theorem 1.3] we see that there is a subgroup $\Gamma_{0}$ such that $\widehat{\Gamma_{0}}$ acts $1$-acylindrically on a profinite tree. Let $H_{0}=\widehat{\Gamma}_{0}\cap H$. As in the proof of Theorem B, by Theorem F, we have an embedding
@@
In the second case we observe first that by Proposition 4.10 $H_{0}\cap G(v)^{g}$ are finitely generated which allows to apply [Zal23a, Theorem 1.4] to deduce that $H_{0}\cap G(v)^{g}$ are free $p pro$ducts of parabolic pro-$p$ groups and a free pro-$p$ group. ∎
@@
Now we are in the position deduce from Proposition 6.4
Theorem C.
  1. (i)
  2. (ii)
    $H$ is a virtually abelian group.
Hence we also have as a particular case
Theorem D.
Let $M$ be a hyperbolic $3$-manifold with cusps and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. Then one of the following statements holds:
  1. (i)
  2. (ii)
    $H$ is a virtually abelian.
@@ 7 The fundamental group of standard compact arithmetic manifolds
It is shown in [BHW11] that standard uniform arithmetic subgroups of $\operatorname{SO}(n,1)$ are virtually compact special and in [PS24, Lemma 6.3 and Theorem 7.4] this was extended to standard non-uniform arithmetic subgroups. Also, it is well-known that the fundamental group of these manifolds is hyperbolic relative to virtually abelian subgroups. Thus, Theorems B and C can be applied to standard arithmetic manifolds.
Theorem 7.2.
Let $M$ be a standard arithmetic manifold and $H$ a finitely generated prosoluble subgroup of the profinite completion of $\pi_{1}(M)$. One of the following holds:
  1. (i)
    $H$ is a subgroup of a free prosoluble product of abelian pro-$p$ groups;
  2. (ii)
    $H$ is abelian.
Moreover, if $\pi_{1}(M)$ is uniform, then $H$ is projective.
8 The fundamental group of compact $3$-manifolds
@@
[WZ18, Theorem 1.3] A finitely generated pro-$p$ subgroup of the profinite completion of the fundamental group of a compact, orientable $3$-manifold $M$ is a free $p pro$duct of pro-$p$ groups from the following list of isomorphism types:
@@
    For $p=2$: in addition to the list of (i) we have $C_{2^{m}}$; $D_{2^{k}}$; $Q_{2^{n}}$; $\mathbb{Z}_{2}\rtimes C_{2}$; the torsion-free extensions of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ by one of $C_{2}$, $C_{4}$, $C_{8}$, $D_{2}$, $D_{4}$, $D_{8}$, $Q_{16}$; the pro-$2$ extension of the Klein-bottle group $\mathbb{Z}\rtimes\mathbb{Z}$; the pro-$2$ completion of all torsion-free extensions of a soluble group $(\mathbb{Z}\rtimes\mathbb{Z})\rtimes\mathbb{Z}$ with a group of order at most $2$.
@@
Let $M$ be a compact orientable $3$-manifold. If $H$ is a finitely generated prosoluble subgroup of $\widehat{\pi_{1}(M)}$, then one of the following statements holds:
@@
    $H$ is isomorphic to a subgroup of a free prosoluble product of pro-$p$ groups from the following list of isomorphism types:
@@
      For $p=2$: in addition to the list of (1) we have $C_{2^{m}}$; $D_{2^{k}}$; $Q_{2^{n}}$; $\mathbb{Z}_{2}\rtimes C_{2}$; the torsion-free extensions of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ by one of $C_{2}$, $C_{4}$, $C_{8}$, $D_{2}$, $D_{4}$, $D_{8}$, $Q_{16}$; the pro-$2$ completion of the Klein-bottle group $\mathbb{Z}\rtimes\mathbb{Z}$; the pro-$2$ completion of all torsion-free extensions of a soluble group $(\mathbb{Z}\rtimes\mathbb{Z})\rtimes\mathbb{Z}$ with a group of order at most $2$;
@@
    $H$ is a subgroup of a central extension of either $D_{2n}$ or tetrahedral group $T$ or octahedral group $O$ by a cyclic group of even order;
@@
    $H$ is a subgroup of the profinite completion of a $3$-dimensional Bieberbach group $B$ (i.e, $B$ is torsion-free virtually $\mathbb{Z}^{3}$);
@@
    $H$ is a subgroup of the profinite completion of a group extension of $\mathbb{Z}^{2}\rtimes_{A}\mathbb{Z}$ by $C$ where $C$ is trivial or $C_{2}$ and $A\in\operatorname{GL}_{2}(\mathbb{Z})$ (and $A$ is an Anosov matrix if $C$ is trivial), or a central extension of $\mathbb{Z}$ by $\mathbb{Z}^{2}$;
@@
    $H$ is an extension of a torsion-free procyclic group $\widehat{\mathbb{Z}}_{\sigma}$ by a subgroup $H_{0}$ of a free prosoluble product of finitely many finite cyclic $p$-groups with $H_{0}$ acting either trivially on $\widehat{\mathbb{Z}}_{\sigma}$ or by inversion.
@@ 8.1 Reduction to the irreducible case
Let $M$ be a compact, orientable, non-geometric $3$-manifold. Since every compact manifold is a retract of a closed manifold, it follows easily that we may reduce to the case in which $M$ is closed.
@@
where each $M_{i}$ is a prime $3$-manifold. A prime $3$-manifold is either irreducible or $S^{2}\times S^{1}$ or the non-orientable $S^{2}$ bundle over $S^{1}$. In the last two cases, the fundamental group is isomorphic to $\mathbb{Z}$ (see [Hem04] and [GSW21, Page 5]).
@@
Let $H$ be a finitely generated prosoluble subgroup of the profinite completion $G=\widehat{\pi_{1}(M)}$ of a reducible $3$-manifold having a non-trivial free product decomposition
@@
where each $M_{i}$ is a irreducible $3$-manifold and $F$ is a free group of finite rank. Then one of the following holds:
@@
    all intersections $H\cap\widehat{\pi_{1}(M_{i})}^{g}$, $g\in G$ are pro-$p$ for some fixed prime $p$ and $\sum_{v}d(H\cap\widehat{\pi_{1}(M_{i})}^{g})\leq d(H)$;
@@ 8.2 Reduction to the geometric cases
Let $M$ be a closed irreducible $3$-manifold. Then the action of $\widehat{\pi_{1}(M)}$ on the standard profinite tree $S$ associated to the JSJ-decomposition of $M$ is $4$-acylindrical.
Then we can apply Theorem F to deduce:
@@
Let $M$ be a closed irreducible $3$-manifold. If $H$ is a finitely generated prosoluble subgroup of $\widehat{\pi_{1}(M)}$, then $H$ embeds into a free profinite product
$\coprod_{v\in V}H_{v}$
of $H$-stabilizers of vertices in $S$ (the standard profinite tree associated to $\widehat{\pi_{1}(M)}$) and one of the following statements holds:
@@
    $H$ is a profinite Frobenius group $\widehat{\mathbb{Z}}_{\pi}\rtimes C$;
@@
If $H$ satisfies $(i)$, then we can apply Theorem 8.1 again and deduce $(i)$ or $(ii)$ of Theorem E. The case $(ii)$ corresponds to Theorem E $(ii)$. Therefore we may assume that $H$ satisfies item $(iii)$ and since $H$ in this case is conjugate into a vertex of $\widehat{\pi_{1}(M)}$ we may assume that $H$ is contained in the profinite completion of a geometric manifold.